OpenStudy (anonymous):

A ball at the end of a 2-m rope swings in a vertical arc. Its speed at the bottom of the swing is 8 m/s. What will be its speed when it is at a position where the rope is horizontal?

OpenStudy (anonymous):

v= square root of (m*length*(1-costheta) ) and theta is 90 right?? :( where am i going wrong?

OpenStudy (jamesj):

Use conservation of energy. At the bottom all the energy is kinetic \[ KE = \frac{1}{2}mv^2 = 32m \] while when the rope is horizontal it has a mixture of KE' (let's call it KE' to differentiate it from the other KE) and PE. The PE is \[ PE = mgh = 19.6m \] as h = 2 Now find the KE' at this point and solve for v' in that KE'

OpenStudy (jamesj):

I.e., KE = PE + KE' You now have expressions for KE and PE, use those to write down an expression for KE' and solve for v.

OpenStudy (anonymous):

I dont understand what your using for mass... ??? 1kg? or 0? If it was 0 then that couldnt be right...

OpenStudy (jamesj):

I didn't put any mass in. I just left that as a variable and that is why 'm' appears in both of the expressions for KE and PE. Keep on going and you'll see you can cancel it out.

OpenStudy (anonymous):

ok we thought it was meters.... i was confused...okokok one second

OpenStudy (anonymous):

-5N

OpenStudy (anonymous):

5 m/s sorry

OpenStudy (jamesj):

right

OpenStudy (anonymous):

velocity at horizontal position=root of(square of(initial velocity)-(2gr)); =root of((8*8)-(2*10*2)) =root of(64-40)=root of 24=5 m/s(about)